# Semantic-guided Prompt Organization for Universal Goal Hijacking against LLMs

### v1

@@ 4 Our Method
In this section, we introduce the three main parts of our simple yet effective universal goal hijacking method: sampling strategy, ranking strategy, and optimization algorithm.
@@ 17                   $I_{f}=I$
19      $\mathcal{BP}$.delete$(I_{f})$$\mathcal{P}$.append$(I_{f})$
@@ 11                   Swap($\mathcal{P}[j]$$\mathcal{P}[j+1])$
4.1 Sampling Strategy
As introduced before, if the training dataset $\mathcal{P}$ of size $N$ contains prompts with similar semantics, it is not conducive for optimizing an adversarial suffix with the “universal” property (see Sec 5.4). Inspired by this, it is intuitive that in order to build a high-quality training dataset, we need to make the prompts in $\mathcal{P}$ to have high semantic diversity.
To be specific, given the big dataset $\mathcal{BP}$ which contains $W$ normal prompts $\{I_{1},I_{2},\dots,I_{W}\}$ and an empty dataset $\mathcal{P}$, a naive method is to find out all the possibility of choosing out $N$ elements from $\mathcal{BP}$ and select the one has the lowest mutual mean semantic similarity to be $\mathcal{P}$. However, from the combination formula $\mathcal{C}(W,N)=\frac{W!}{N!(W-N)!}$, it is obvious that the time and resource consumption is unacceptable when $W$ is big. Thus our sampling strategy is based on the greedy algorithm and aims to find an approximate solution. Specifically, the sampling strategy contains three steps. ❶ Calculate the semantic similarity between all the pairs in dataset $\mathcal{BP}$ and add the pair that has the lowest similarity to the training dataset $\mathcal{P}$. ❷ Select a prompt $\hat{I}$ from $\mathcal{BP}$ which has the accumulative total lowest semantic similarity with all existing prompts in $\mathcal{P}$ and add the prompt $\hat{I}$ into $\mathcal{P}$. ❸ Repeat the second step until the number of prompts in training dataset $\mathcal{P}$ is $N$. We demonstrate the sampling strategy in Algorithm 1. From line 1 to 1, there shows the details of step ❶. From line 1 to 1, there shows the procedures of step ❷. For the similarity evaluation metric, we conduct experiment and find cosine similarity as a good choice.
@@ 4.2 Ranking Strategy
With the sampled training dataset $\mathcal{P}$ from Sec. 4.1, our ranking strategy is target response-related and contains two steps. ❶ Calculate the similarity between prompts in $\mathcal{P}$ and target response $R^{T}$, then save the similarity into list $\mathcal{Q}$. ❷ Sort the prompts in $\mathcal{P}$ with the sort of $\mathcal{Q}$. We demonstrate the ranking strategy in Algorithm 2. From line 2 to 2, there shows the details of step ❶. From line 2 to 2, there shows the procedures of step ❷. For the similarity evaluation metric, we also use cosine similarity. Through the experiment, we find sorting $\mathcal{Q}$ with descending order can successfully lead to a faster convergence speed of optimization procedure than random sort. That is, we suggest putting the prompt that has the highest semantic similarity with the target response into the optimization algorithm first and followed by prompts whose semantic similarity with the target response gradually decreases. Furthermore, we try sorting $\mathcal{Q}$ with ascending order and find the convergence speed is even faster than descending order. However, this usually leads to a relatively low ASR (about 10% lower than descending order). Since for the universal goal hijacking task, ASR is the dominant metric, we choose descending order as the default choice in our experiment section. To summarize, no matter the order, we find ranking is better than random sorting for the universal goal hijacking task.
4.3 Optimization Algorithm
To solve the optimization task Eq. 5, recent adversarial attack methods which work on discrete tokens are significant references (e.g., GCG [36]). However, GCG concentrates on jailbreak attacks in which the whole malicious input prompts are determined by the adversary, in contrast to the goal hijacking attack in which the user’s normal prompt is unknown. Thus our algorithm only follows the idea of GCG in processing discrete tokens while different in input and loss design.
With regards to the algorithm, the optimization is on the loss which only aggregates part of the training dataset in early iterations and gradually increases the number of input prompts until the number is the same as the size of the training dataset. We call this algorithm the I-UGH. Here we demonstrate I-UGH in Algorithm 3.
In line 3, initialize the number of prompts ($n_{c}$) that participated in loss gradient calculation. In line 3, add fixed suffix $S_{1:q}$ to $n_{c}$ input prompts of the training dataset to calculate the loss and gradient. In line 3, compute the top-k token substitutions from the vocabulary as the candidate replacements $\mathcal{S}_{i}$ for each token position of $S_{1:q}$. From line 3 to 3, compute the candidate set $\widetilde{S}_{1:q}$ for all tokens in suffix and randomly select $B$ tokens. In line 3, evaluate the candidate in $\widetilde{S}_{1:q}$ and select the one that achieves the smallest loss as the replacement. In line 3, $n_{c}$ needs to increase when the suffix can achieve high ASR on the part training dataset $\mathcal{P}_{1:n_{c}}$. Here to avoid overfitting of the suffix, we only require the suffix to succeed on most part of the prompts and use a threshold (0.8 in our experiment) to control this. If ASR is higher than the threshold, then increasing the $n_{c}$.
@@ 5.1 Experimental Setups
Datasets and models. In our evaluations, we use the normal prompts collected from the Alpaca dataset [40] to construct the training dataset and test dataset. Alpaca is a popular public prompt dataset open-sourced by Stanford with about 100,000 downloads per month. We utilized Llama-2-7b-chat-hf [41], Vicuna-7b-v1.5 [42] and Guanaco-7B-HF [43], Mistral-7B-Instruct [44] as the victim models. These models are classical open-source models that are popular on the Hugging Face platform [45].
Implementation details of our method. For the big normal prompt dataset $\mathcal{BP}$ and training dataset $\mathcal{P}$, the size is 1,000 and 50. For the test dataset $\mathcal{P}_{test}$, the size is 1,000. We refer to the setting of GCG and set the hyperparameters of our method as follows: the batch size $B$ is 128, the top-k value is 64, the fixed total iteration number $T$ is 1,000 and the suffix length $q$ is 128. The semantic extraction function $\Theta(\cdot)$ is realized by extracting the embedding of the last hidden state in LLM. All the experiments were run on an Ubuntu system with an NVIDIA A100 Tensor Core GPU of 80G RAM.
Input: Initial suffix $S_{1:q}$, Training prompt dataset $\mathcal{P}$ of size $N$, Batch size $B$, Iterations $T$, Loss $\mathcal{L}$, LLM model $\mathbf{LM}(\cdot)$
Output: Optimized suffix $S_{1:q}$
1 $n_{c}$ = 1
2 for $t=1\ \mathrm{to}\ {T}$ do
3       for $i=1\ \mathrm{to}\ {q}$  do
[…]
5             $G_{t}\leftarrow\sum_{j=1}^{n_{c}}-\bigtriangledown_{e_{S_{i}}}\mathcal{L}(I_{j},S_{1:q},\mathbf{LM})$
[…]
7             $\mathcal{S}_{i}\leftarrow$Topk$(G_{t})$
8      for $b=1\ \mathrm{to}\ {B}$  do
[…]
10             $\widetilde{S}_{1:q}^{(b)}\leftarrow S_{1:q}$
[…]
16       if $S_{1:q}$ succeeds on $\mathcal{P}_{1:n_{c}}$  then
17             if $n_{c}<N$ then
18                   $n_{c}\leftarrow n_{c}+1$
19             else
20                   return $S_{1:q}$
21
22
Algorithm 3 I-UGH
Baselines. We consider M-GCG [17], a newly proposed state-of-the-art method that incorporates the concept of momentum into the GCG and efficiently achieves high ASR. It outperforms most goal hijacking methods such as [46, 47] and is the only work for the universal goal hijacking task.
@@ 18                   $n_{c}\leftarrow n_{c}+1$
Evaluation protocols and metrics. To evaluate the effectiveness of the method across different target responses, we design target responses from 10 malicious categories (threatening, bomb, fraud, virus, murder, phishing, financial, drug, racism, and suicide, listed in Table IV of Appendix). The categories are summarized from the famous dataset AdvBench [48]. We evaluate the algorithm from two aspects: attack success rate and time consumption. For the metric of time consumption, it is not suitable to use time such as hour or minute since different GPU servers may lead to distinct results. Thus we evaluate the time consumption of each part in I-UGH (Algorithm 3). For each iteration, there are four parts: calculate gradient (line 3-3), select candidate (line 3-3), calculate the best suffix (line 3), check results (line 3-3). In Table I, we evaluate the time consumption when $n_{c}$ are 1 and 50. We can find that the “calculate best” part takes most of the time when $n_{c}$ = 1, accounting for about 63%. When $n_{c}$ = 50, excluding the “select candidate” part, the time consumption of other parts significantly increases. We list the magnification of $n_{c}$ = 1 and 50 columns in the “scale” column. Particularly, the time consumption of the “calculate best” and “check result” parts takes about 50$\times$ magnification. They account for about 76% and 20% respectively, a total of 96%. Since they are proportional to $n_{c}$ and take a huge account of time consumption among the four parts, thus we use the number of accumulation of $n_{c}$ (#NC) in all the iterations as the metric for time consumption.
TABLE II: Comparison with baseline on llama-2.
@@
ASR (%) $\uparrow$ M-GCG (1000) 65.1 4.3 95.6 89.7 98.8 97.8 98.6 93.5 99.2 95.8 83.84
M-GCG (500) 24.8 0.0 79.8 0.0 93.6 88.6 94.3 0.0 92.8 68.7 54.26
POUGH (ours) 92.6 93.5 85.7 97.3 92.0 85.0 88.9 82.9 98.7 92.8 90.94
Time (#NC) $\downarrow$ M-GCG (1000) 27050 50000 50000 50000 50000 50000 50000 50000 50000 50000 47705.0
M-GCG (500) 25000 25000 25000 25000 25000 25000 25000 25000 25000 25000 25000.0
POUGH (ours) 2092 23478 1991 2196 9864 6110 5870 2589 6109 3528 6382.7
@@ 5.2 Main Results
Compare with baseline. In Table II, we show the ASR and time comparison between M-GCG and our method. For the M-GCG method, we perform it for 1,000 iterations. If the method does not converge in 1,000 iterations, we will stop the method and take the last suffix for testing. The results are shown in the “M-GCG (1000)” row. We also take the suffix at the 500-th iteration for testing, shown in the “M-GCG (500)” row. For the time consumption of “M-GCG (1000)”, since in each iteration, they take all the prompts (50) in $\mathcal{P}$ for gradient calculation, then the #NC of them is #iteration $\times$ 50 and 50,000 if converge/not converge. This is the same for the “M-GCG (500)” row.
From the Table, comparing with the “M-GCG (1000)” row, we can find that our method achieves better ASRs than M-GCG (90.34% vs. 83.84%) while being much more efficient than it (only using 13.7% time). Compared with the “M-GCG (500)”, the advantage of our method is more obvious due to the bad performance of M-GCG (54.26% ASR).
Performance on different models. In Table III, we show the performance of our method on more different target LLMs, including vicuna, mistral, and guanaco. We can find that our method can hijack all LLMs efficiently and effectively. On average, the method can achieve high ASR (more than 80%). Also, we can find that optimization time on mistral is obviously higher than that on vicuna and guanaco, which reflects the mistral model is harder for universal goal hijacking.

### v2

@@ 4 Our Method
The algorithms are introduced below, with their complexity analyses provided in Appendix 10.
Input: Initial suffix $S_{1:q}$, Training prompt dataset $\mathcal{P}$ of size $N$, Target response $R^{T}$, Batch size $B$, Iterations $T$, LLM model $\mathbf{LM}(\cdot)$
Output: Optimized suffix $S_{1:q}$
1 $n_{c}$ = 1
2 for $t=1\ \mathrm{to}\ {T}$ do
3       for $i=1\ \mathrm{to}\ {q}$  do
[…]
5             $G_{t}\leftarrow-\bigtriangledown_{e_{S_{i}}}\sum_{I\in\mathcal{P}_{1:n_{c}}}\mathcal{L}(I,S_{1:q},R^{T},\mathbf{LM})$
[…]
7             $\mathcal{V}_{i}\leftarrow$Topk$(G_{t})$
8      for $b=1\ \mathrm{to}\ {B}$  do
[…]
10             $\widetilde{S}_{1:q}^{(b)}\leftarrow S_{1:q}$
[…]
16       if $S_{1:q}$ succeeds on $\mathcal{P}_{1:n_{c}}$  then
17             if $n_{c}<N$ then
18                   $n_{c}\leftarrow n_{c}+1$
19             else
20                   return $S_{1:q}$
21
22
Algorithm 1 I-UGH
4.1 Optimization Algorithm
To design an efficient optimization algorithm, we first observe the prompt-specific suffix that is generated for a single user prompt. Through experiment, we find although the prompt-specific suffix does not satisfy the requirement of universal goal hijacking (i.e., has high ASR across different user prompts), the ASR is not zero (0.84% in Table II), which means it has weak universality. Meanwhile, the generation speed of the prompt-specific suffix is fast. Then comes an intuitive idea: we can generate the prompt-specific suffix first with a single user prompt and gradually optimize it to have “universality”. To be specific, similar to the state-of-the-art universal goal hijacking method M-GCG [7], our algorithm also follows the optimization idea of GCG since it works well on optimizing discrete tokens. The difference is that we suggest starting with only one prompt as input and gradually increasing the number of prompts that participated in loss calculation until it matches the size of the training dataset. The optimization algorithm is much more efficient than the M-GCG. We demonstrate it in Algorithm 1.
[Line 1] The algorithm starts by setting the number of prompts ($n_{c}$) that participated in loss calculation to be 1. [Line 1 to 1] (Get token substitutions for each token in suffix $S_{1:q}$) In line 1, use Eq. 4 to calculate the sum of losses for $n_{c}$ user prompts and calculate the gradient $G_{t}$ for the token $S_{i}$ in the suffix $S_{1:q}$. In line 1, with gradient $G_{t}$ for token $S_{i}$, select the top-k token substitutions from the vocabulary to be $\mathcal{V}_{i}$. [Line 1 to 1] (Build suffix candidate set $\widetilde{S}_{1:q}$ of size $B$) In line 1, initialize a candidate suffix $\widetilde{S}_{1:q}^{(b)}$ to be same as $S_{1:q}$ first. Then, in line 1, replace the token $S_{i}$ in candidate suffix $\widetilde{S}_{1:q}^{(b)}$ according to token substitutions $\mathcal{V}_{i}$. Each candidate suffix $\widetilde{S}_{1:q}^{(b)}$ in the set $\widetilde{S}_{1:q}$ just has one token difference with suffix $S_{1:q}$. [Line 1] (Select a suffix from the suffix candidate set) For each candidate suffix in $\widetilde{S}_{1:q}$, use Eq. 4 to calculate the sum of losses for $n_{c}$ user prompts and select the one that achieves the smallest loss. It is the new suffix for further optimization in the next iteration. [Line 1 to 1] (Gradually increase prompts participated in optimization) $n_{c}$ needs to increase when the new suffix can achieve high ASR on the part of the training dataset (i.e., $\mathcal{P}_{1:n_{c}}$). To avoid overfitting, we only require the suffix to succeed on most parts of the prompts and use a threshold (0.8 in our experiment) to control this. If ASR is higher than the threshold, then increasing the $n_{c}$.
@@ 17                   $\hat{I}=I$
19      $\mathcal{BP}$.delete$(\hat{I})$$\mathcal{P}$.append$(\hat{I})$
@@ 20       $n_{c}$leftarrow $n_{c}$ + 1
4.2 Sampling Strategy
Existing attack methods [30, 7] put too much emphasis on the algorithm design. However, for the universal goal hijacking, we propose prompt is a crucial factor in cooperation with the algorithm that cannot be ignored.
Note the optimization algorithm needs to deal with a large volume of prompts, which leads to high computational intensity, it is obvious that a training dataset with a small size and high quality is preferred. Inspired by this, we propose the problem setting should be extended a bit. That is, we can collect a lot of normal user prompts $\mathcal{BP}$ from the web and select a small subset $\mathcal{P}$ of high quality from it to be the training dataset. The idea behind the sampling strategy comes from the following observations. If the prompts in $\mathcal{P}$ all have similar semantics as the prompt “Provide three pieces of advice for maintaining good health.”, even the adversarial suffix can achieve the 100% ASR on $\mathcal{P}$, the universality of the suffix is low on the test dataset (5% ASR, verified in Sec. 5.4). Inspired by this, we suggest constructing the dataset $\mathcal{P}$ with high semantic diversity and we also find that the selection of prompts can be irrelevant to the target response $R^{T}$.
To be specific, given the big dataset $\mathcal{BP}$ which contains $W$ normal prompts and an empty dataset $\mathcal{P}$, a naive method is to find out all the possibilities of choosing out $N$ elements from $\mathcal{BP}$ and select the one has the lowest mutual mean semantic similarity to be $\mathcal{P}$. However, from the combination formula $\mathcal{C}(W,N)=\frac{W!}{N!(W-N)!}$, it is obvious that the time and resource consumption is unacceptable when $W$ is big. Thus our sampling strategy is based on the greedy algorithm and aims to find an approximate solution. Specifically, the sampling strategy contains three steps. ❶ Calculate the semantic similarity between all the pairs in dataset $\mathcal{BP}$ and add the pair that has the lowest similarity to the training dataset $\mathcal{P}$. ❷ Select a prompt $\hat{I}$ from $\mathcal{BP}$ which has the accumulative total lowest semantic similarity with all existing prompts in $\mathcal{P}$ and add the prompt $\hat{I}$ into $\mathcal{P}$. ❸ Repeat the second step until the number of prompts in training dataset $\mathcal{P}$ is $N$. We demonstrate the sampling strategy in Algorithm 2. From line 2 to 2, there shows the details of step ❶. From line 2 to 2, there shows the procedures of step ❷. For the similarity evaluation metric, we find cosine similarity as a good choice.
@@ 4.3 Ranking Strategy
Since our optimization algorithm gradually increases the number of prompts participating in loss calculation, will different sequences of prompts lead to distinct convergence speeds? The answer is YES (verified in Sec. 5.3). There comes a question that how to define the priority of the prompts?
For this question, given the training dataset $\mathcal{P}$ from Sec. 4.2, the goal is to rank the prompts in $\mathcal{P}$ and achieve the adversarial suffix $S$ efficiently. We find the target response can provide guidance on the ranking. That is, we can use the semantic similarity between each prompt and the target response as a metric. Inspired by this idea, our ranking strategy is target response-related and contains two steps. ❶ Calculate the similarity between prompts in $\mathcal{P}$ and target response $R^{T}$, then save the similarity into list $\mathcal{Q}$. ❷ Sort the prompts in $\mathcal{P}$ with the sort of $\mathcal{Q}$. We demonstrate the ranking strategy in Algorithm 3. For the similarity evaluation metric, we also use cosine similarity. Through the experiment, we find sorting $\mathcal{Q}$ with descending order can successfully lead to a faster convergence speed of optimization procedure than random sort. That is, we suggest putting the prompt that has the highest semantic similarity with the target response into the optimization algorithm first and followed by prompts whose semantic similarity with the target response gradually decreases. Note that semantic similarity may not be the best criterion for ranking but it is better than random sort.
@@ 5.1 Experimental Setups
Datasets and models. In our evaluations, we use the normal user prompts (easy to achieve on the web) collected from the Alpaca dataset [35] to construct the training dataset and test dataset. Alpaca is a popular public dataset open-sourced by Stanford which contains diverse prompts and achieves about 100,000 downloads per month. We utilized Llama-2-7b-chat-hf [36], Vicuna-7b-v1.5 [37] and Guanaco-7B-HF [38], Mistral-7B-Instruct [39] as the victim models. These models are classical open-source models that are popular on the Hugging Face platform [40].
Implementation details of our method. For the big normal prompt dataset $\mathcal{BP}$ and training dataset $\mathcal{P}$, the size is 1,000 and 50. For the test dataset $\mathcal{P}_{test}$, the size is 1,000. The hyperparameters of our method are as follows: the batch size $B$ is 128, the top-k value is 64, the fixed total iteration number $T$ is 1,000 and the suffix length $q$ is 128. The semantics extraction function $\Theta(\cdot)$ is realized by extracting the embedding of the last hidden state in LLM [41]. All the experiments were run on an Ubuntu system with an NVIDIA A100 Tensor Core GPU of 80G RAM.
Baselines. We choose classical and popular optimization algorithms for discrete tokens (i.e., GCG [30], MAC [42], AutoDAN [43], and AmpleGCG [44]) and adapt them to fit the setting of goal hijacking to see their performance. For the universal goal hijacking task, we choose classical and popular handcrafted methods HouYi [4] and TensorTrust [5] as well as the gradient-based optimization method M-GCG [7]. Note that to our best knowledge, M-GCG is the first and only optimization method designed for the universal goal hijacking task and achieves the best ASR. For all the baselines and our method, the upper limit for #NC is 25,000 since it needs more than one day on A100 GPU, which is a long time.
@@
TABLE II: Comparison with baselines on llama-2.
@@
ASR (%) $\uparrow$ GCG-hijacking 0.2 0.0 0.1 7.1 0.2 0.0 0.5 0.0 0.3 0.0 0.84
MAC-hijacking 0.0 0.0 0.7 0.2 0.2 0.8 0.0 0.1 1.6 0.0 0.36
AutoDAN-hijacking 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.00
AmpleGCG-hijacking 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.00
[…]
TensorTrust 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.00
M-GCG 24.8 0.0 79.8 0.0 93.6 88.6 94.3 0.0 92.8 68.7 54.26
I-UGH (ours) 92.5 84.2 86.3 82.8 88.7 88.9 70.8 79.7 88.8 91.9 85.50
POUGH (ours) 92.6 93.5 94.4 97.3 92.8 92.0 97.1 82.9 98.7 92.8 93.41
Time (#NC) $\downarrow$ GCG-hijacking 303 567 267 435 450 205 578 909 445 461 462.0
MAC-hijacking 546 490 362 479 344 309 758 299 518 600 470.5
AutoDAN-hijacking 500 500 500 500 500 500 500 500 500 500 500.0
AmpleGCG-hijacking - - - - - - - - - - -
[…]
TensorTrust - - - - - - - - - - -
M-GCG 25000 25000 25000 25000 25000 25000 25000 25000 25000 25000 25000.0
I-UGH (ours) 11280 4672 2844 7294 11722 2795 19444 9726 2340 4014 7613.1
POUGH (ours) 2092 23478 2306 6105 3049 3406 3600 2589 6109 3528 5626.2
@@
Evaluation protocols and metrics. To evaluate the effectiveness of the method across different target responses, we design target responses from 10 malicious categories (threatening, bomb, fraud, virus, murder, phishing, financial, drug, racism, and suicide, listed in the Appendix 11). The categories are summarized from the famous dataset AdvBench [45]. We evaluate the algorithm from two aspects: attack success rate and time consumption. For the metric of time consumption, it is not suitable to use time such as hours or minutes since different GPU servers may lead to distinct results. Thus we evaluate the time consumption of each part in our optimization algorithm (Algorithm 1). For each iteration, there are four parts: calculate gradient (line 1-1), select candidate (line 1-1), calculate the best suffix (line 1), check results (line 1-1). In Table I, we evaluate the time consumption when $n_{c}$ are 1 and 50. We can find that the “calculate best” part takes most of the time when $n_{c}$ = 1, accounting for about 63%. When $n_{c}$ = 50, excluding the “select candidate” part, the time consumption of other parts significantly increases. We list the magnification of $n_{c}$ = 1 and 50 columns in the “scale” column. Particularly, the time consumption of the “calculate best” and “check result” parts takes about 50$\times$ magnification. They account for about 76% and 20% respectively, a total of 96%. Since they are proportional to $n_{c}$ and take a huge account of time consumption among the four parts, thus we use the number of accumulation of $n_{c}$ (#NC) in all the iterations as the metric.
@@ 5.2 Main Results
Compare with baseline. In Table II, we show the ASR and time comparison between baselines and our method. The baselines GCG-hijacking, MAC-hijacking, AutoDAN-hijacking and AmpleGCG-hijacking exploit the popular optimization algorithm GCG [30], MAC [42], and AutoDAN [43], AmpleGCG [44] respectively while adapting them to goal hijacking task. Note that without making major modifications to the algorithm, they are only able to generate a prompt-specific suffix. We list them here to show the bad “universality” of suffixes generated for the typical goal hijacking task. Since AmpleGCG is a generative-based method, thus its time consumption is labeled as “-”. With regard to methods for universal goal hijacking methods, we show the performance of HouYi [4], TensorTrust [5] and M-GCG [7]. For HouYi [4] and TensorTrust [5], since they are handcrafted, the time consumption is labeled as “-”. With regard to our method, we show the performance of the I-UGH algorithm and the POUGH method (i.e., I-UGH combined with two prompt organization strategies).
From the Table, we can find that the ASRs of GCG-hijacking, MAC-hijacking, and AutoDAN-hijacking are near zero, which means the optimization algorithms for jailbreaking, even modified to adapt to the goal hijacking task, can not achieve good results on the universal goal hijacking task. Also, the result reflects that the prompt-specific suffix has weak universality since the ASR is not zero. With regard to methods designed for the universal goal hijacking task, the handcrafted methods HouYi and TensorTrust achieve bad attack performance (ASRs less than 1%). Due to the extreme inefficiency of M-GCG, which requires gradient calculations for all training prompts in each iteration, we had to impose a practical time constraint of approximately one day on an A100 GPU for benchmarking. M-GCG achieves a higher attack performance (higher than 50%). Compared with the M-GCG, our method (I-UGH) achieves higher ASRs than M-GCG (85.50% vs. 54.26%) while being much more efficient (only using 26.2% time). Furthermore, our proposed POUGH method achieves the highest ASRs (93.41%) with nearly a fifth of the time consumption compared with M-GCG.
Performance on different models. In Table III, we show the performance of our method on more target LLMs, including vicuna, mistral, and guanaco. We can find that our method can hijack all LLMs efficiently and effectively. On average, the method can achieve high ASR (more than 80%). Also, we can find that optimization time on mistral is obviously higher than that on vicuna and guanaco, which reflects the mistral model is harder for universal goal hijacking.